Car A leaves Toronto for Montreal, 500km away, at an average speed of 80km/h. Car B leaves Montreal for Toronto on the same highway 2h later at 100 km/h. How far are they from Toronto when they pass?
equation for this please.. tnx in advance. =D

Answer by stanbon(75887)

Car A leaves Toronto for Montreal, 500km away, at an average speed of 80km/h. Car B leaves Montreal for Toronto on the same highway 2h later at 100 km/h. How far are they from Toronto when they pass?
---------
Car A DATA:
Rate = 80 kph ; Distance = x km ; Time = d/r = x/80 hrs
-------------
Car B DATA:
Rate = 100 kph ; Distance = (500-x) km ; time = d/r = (500-x)/100 hrs
--------------
EQUATION:
Car A time - Car B time = 2 hrs
x/80 - (500-x)/100 = 2
Multiply thru by 800 to get:
10x - 8(500-x) = 1600
10x - 4000 + 8x = 1600
18x = 5600
x = 311.1111 km (Distance from Toronto when they pass)
==============
Cheers,
Stan H.